I tried to find the answer to this question, but was totally confused.

One card is selected at random from a standard 52-card deck of playing cards. Find the probability that the card selected is not a spade.

Found 2 solutions by checkley71, jm:
Answer by checkley71(8403)

Non spades=52-13=39 cards.
success/total chances=39/52=3/4=.75=75% chance of not selecting a spade.

Answer by jm(1)


1/4 = 3/4 = 0.75 = 75%